# Safe Model-Based Multi-Agent Mean-Field Reinforcement Learning

### v1

@@ 4 Safe-$\text{M}^{3}$-UCRL
We summarize the model-based learning protocol used by Safe-$\text{M}^{3}$-UCRL in Algorithm 1. The first step consists of computing constants $C_{n,t}$ (see Appendix A) introduced in Lemma 1. The second step is optimizing the objective in Section 4. The third and fourth steps collect trajectories from the representative agent and update the calibrated model. While the learning protocol is model-based, the subroutine in Line 3 can be implemented with either model-based or model-free algorithms proposed for the Mean-Field Control due to our reformulation in Section 4. In Section C.3, we introduce modifications of well-known algorithms for optimizing the mean-field setting.
5 Experiments
In this section, we demonstrate the performance of Safe-$\text{M}^{3}$-UCRL on the vehicle repositioning problem of a ride-hailing platform. Since ride-hailing services, such as Uber and Lyft, gained popularity and market share, vehicle repositioning has been a long-standing challenge for these platforms, i.e., moving idle vehicles to areas with high-demand potential. In the competitive environment, the operator significantly increases profit by successfully repositioning idle vehicles to high-demand areas. Nevertheless, there might exist regulations imposed by the country or city that enforce service providers to either guarantee fair service accessibility or restrict the number of vehicles in districts with high traffic density. Such restrictions prevent operators from greedily maximizing the profit and can be encapsulated by some dispersion metric such as entropy. Solving this problem helps prevent prolonged vehicle cruising and extensive passenger waiting times in the demand hotspots, increasing the service provider’s efficiency and reducing its carbon footprint. Existing approaches to vehicle repositioning range from static optimization over a queuing network [56, 7], model predictive control [27], to reinforcement learning [54, 35, 37]. The main advantage of Safe-$\text{M}^{3}$-UCRL is the capability of controlling a large fleet of homogeneous vehicles and enforcing efficient coordination to match the spatiotemporal demand distribution. Additionally, the safety constraint introduced into the model guarantees service accessibility by ensuring idle vehicles are spreading over the study region. Although accessibility has not been widely discussed in the literature on vehicle repositioning, it is expected to be an important fairness constraint when shared mobility services become a prevailing travel mode [48].
Observed demand used as a target distribution $\rho_{0}$
$\text{M}^{3}$-UCRL with known dynamics
Safe-$\text{M}^{3}$-UCRL with known dynamics
Safe-$\text{M}^{3}$-UCRL with unknown dynamics
Figure 2: Safe-$\text{M}^{3}$-UCRL performance in Shenzhen. We show the final mean-field distributions $\mu_{T}$ for $T=12$ and the entropic safety constraint $h_{C}(\cdot)$ (Section 3.3). We set the safety threshold as $C=0.85\log(k^{2})$ with $k=25$ which represents 85% of the maximum Shannon entropy. The observed demand distribution $\rho_{0}$ in (a) horizontally spreads across the city center, which amounts to 67% of the maximum entropy, while residential areas in the northwest and especially northeast show only glimpses of demand. Still, the service providers might be obligated to guarantee coverage in these areas. (b) shows that $\text{M}^{3}$-UCRL , the benchmark model without safety constraint, closely matches the target distribution, resulting in safety constraint violation as shown in Figure 2(b). (c) shows that our model with known dynamics achieves coverage in all residential areas while keeping the majority of taxis in demand hotspots. (d) shows the competitive performance of our model even when faced with unknown dynamics.
Modeling. Ride-hailing operations can be modeled as sequential decision-making, which consists of passenger trips followed by repositioning trips operated by a central controller as illustrated in Figure 1. We assume that the controller has access to the locations of taxis in its fleet and that real-time communication of repositioning actions to the taxi drivers is possible via electronic devices. Nevertheless, since the fleet is operating in a noisy traffic environment, repositioning usually cannot be executed perfectly. We assume that taxis can move freely within the area of our interest, which is represented by a two-dimensional unit square, i.e., the state-space $\mathcal{S}=[0,1]^{2}$, and repositioning actions are taken from $\mathcal{A}=[-1,1]^{2}$. The objective of our model is to satisfy the demand in the central district of Shenzhen while providing service accessibility in the wider city center. We restrict our modeling horizon to three evening peak hours which are discretized in fifteen-minute operational intervals, i.e., $T=12$, and each episode $n$ represents one day (for more details, see Appendix D). We implicitly model the profit by the negative of the Kullback-Leibler divergence between the taxis’ distribution $\mu_{n,t}$ and demand for taxi service denoted as $\rho_{0}$, i.e., $r(s_{n,t},\mu_{n,t},a_{n,t})=-D_{KL}(\rho_{0}||\mu_{n,t})$. In particular, the demand distribution $\rho_{0}\in\mathcal{P}(\mathcal{S})$ represents a probability of a trip originating in the infinitesimal neighborhood of state $s\in\mathcal{S}$ during peak hours (see Figure 1(a)). We estimate the demand distribution $\rho_{0}$ from the taxi trajectories collected in Shenzhen, China, in 2016 and keep it fixed for every step $t$. If the passenger’s trip originates at $s\in\mathcal{S}$ the likelihood of its destinations is defined by the mapping $\Phi:\mathcal{S}\rightarrow\mathcal{P}(\mathcal{S})$ which we fit from the trip trajectories (see Appendix D). We use $\Phi(\cdot)$ to define sequential transitions by first executing passenger trips followed by vehicle repositioning. Formally, the unknown system dynamics are given by $f(s_{n,t},\mu_{n,t},a_{n,t})=\text{clip}(s^{\Phi}_{n,t}+a_{n,t},0,1)+\varepsilon_{n,t}$, where $s^{\Phi}_{n,t}=\text{sample}(\Phi(s_{n,t}))$ and $\varepsilon_{n,t}\sim\text{TN}(0,\sigma^{2}I_{2})$ is a Gaussian with a known variance $\sigma^{2}$ truncated at the borders of $\mathcal{S}$ and $I_{2}$ is the $2\times 2$ unit matrix. We use entropic safety constraint $h_{C}(\mu_{n,t})=-\int_{\mathcal{S}}\log(\mu_{n,t}(s))\mu_{n,t}(ds)-C$ to enforce the service accessibility across all residential areas (see Figures 1(c) to 1(d)). Therefore, the optimization objective in Section 4 trades off between satisfying the demand $\rho_{0}$ and adhering to accessibility constraint imposed by $h_{C}(\cdot)$.
@@ 5 Experiments
We use neural network to parametrize the policy profile $\bm{\pi}_{n}=(\pi_{n,0},\ldots,\pi_{n,T-1})$, for every episode $n$, during the optimization of Section 4. The optimization is done by a mean-field back-propagation through time (MF-BPTT) (see Section C.3.1). In our experiments, a single neural network shows enough predictive power to represent the whole policy profile, but using $T$ networks, one for each individual policy $\pi_{n,t}$ is a natural extension. We use a Probabilistic Ensemble to represent a statistical model of system dynamics, which is elaborated in Section C.1. We represent the mean-field distribution by discretizing the state space into $k\times k$ grid with $k=25$ and assigning the probability of the representative agent residing within each cell. The initial mean-field distributions $\mu_{n,0}$ follow the uniform distribution to maximize the Shannon entropy, ensuring the safety at $t=0$, i.e., $h_{C}(\mu_{n,0})\geq 0$ for every episode $n$. We set the safety threshold $C$ as a percentage $p$ of the maximum Shannon entropy, i.e., $C=p\log(k^{2})$.
(a) Objective from Equation 8a achieved
over 200 training episodes
(b) Constraint violations comparison
against the benchmark
Figure 3: Safe-$\text{M}^{3}$-UCRL training and evaluation. We train 10 randomly initialized policy profiles and show that: In (a) the performance of policy profiles during the exploration is quite volatile, but around episode 80, policy profiles become confident about the underlying dynamics and quickly converge to a solution as if the dynamics were known; In (b) our policy profiles deployed in the true environment never violate the entropic safety constraint $h_{C}(\cdot)\geq 0$ for $C=0.85\log(k^{2})$ with $k=25$, while the benchmark model $\text{M}^{3}$-UCRL trained without the safety constraint starts displaying unsafe behavior after only two steps.

### v2

@@ 3.3. Statistical Model and Safety Implications
4. Safe-M3-UCRL
In this section, we introduce a model-based approach for the Safe Mean-Field Reinforcement Learning problem that combines the safety guarantees in Corollary 2 with upper-bound confidence interval optimization. At the beginning of each episode $n$, the representative agent constructs the confidence set of transitions $\mathcal{F}_{n-1}$ (see Equation 6) given the calibrated statistical model and previously observed data and selects a safe optimistic policy profile $\bm{\pi}_{n}^{*}$ to obtain the highest value function within $\mathcal{F}_{n-1}$ while satisfying the safety constraint derived in Corollary 2. In particular, the optimal policy profile $\bm{\pi}^{*}$ from Section 3.1 is approximated at the episode $n$ by
$\bm{\pi}_{n}^{*}{}={}\operatorname*{arg\,max}_{\bm{\pi}_{n}\in\Pi}{}$ $\max_{\tilde{f}_{n-1}\in\mathcal{F}_{n-1}}\mathbb{E}\Bigg{[}\sum_{t=0}^{T-1}r(\tilde{z}_{n,t})\Big{\rvert}\tilde{\mu}_{n,0}{}={}\mu_{0}\Bigg{]}$ (7a)
subject $\text{subject to}\quad\tilde{a}_{n,t}$ ${}={}\pi_{n,t}(\thinspace_{n,t},\tilde{\mu}_{n,t})$ (7b)
$\thinspace{}_{n,t{}+{}1}$ ${}={}\tilde{f}_{n-1}(\tilde{z}_{n,t}){}+{}\varepsilon_{n,t}$ (7c)
$\tilde{\mu}_{n,t{}+{}1}$ ${}={}U(\tilde{\mu}_{n,t},\pi_{n,t},\tilde{f}_{n-1})$ (7d)
$h_{C}(\tilde{\mu}_{n,t{}+{}1})$ ${}\geq{}L_{h}C_{n,t{}+{}1},$ (7e)
with $\tilde{z}_{n,t}{}={}(\thinspace_{n,t},\tilde{\mu}_{n,t},\tilde{a}_{n,t})$. Section 4 optimizes over the function space $\mathcal{F}_{n-1}$ which is usually intractable even in bandit settings Dani et al. (2008). Additionally, it must comply with the safety constraint Equation 7e, further complicating the optimization. We utilize the hallucinated control reparametrization and the log-barrier method to alleviate these issues. After the reformulation of the problem, model-free or model-based mean-field optimization algorithms can be applied to find policy profile $\bm{\pi}_{n}^{*}$ at the beginning of episode $n$.
We use an established approach known as Hallucinated Upper Confidence Reinforcement Learning (H-UCRL) Moldovan et al. (2015); Curi et al. (2020); Pásztor et al. (2023) and introduce an auxiliary function $\eta:\mathcal{Z}\rightarrow[-1,1]^{p}$, where $p$ is the dimensionality of the state space $\mathcal{S}$, to define hallucinated transitions
Notice that $\tilde{f}_{n-1}$ is calibrated for any $\eta(\cdot)$ under Assumption 4, i.e., $\tilde{f}_{n-1}\in\mathcal{F}_{n-1}$. Assumption 4 further guarantees that every function $\tilde{f}_{n-1}$ can be expressed in the auxiliary form Equation 8
$\begin{split}&\forall\tilde{f}_{n-1}\in\mathcal{F}_{n-1}\;\exists\eta:\mathcal{Z}\rightarrow[-1,1]^{p}\;\text{such that}\\ {}&\tilde{f}_{n-1}(z){}={}\bm{m}_{n-1}(z){}+{}\beta_{n-1}\bm{\Sigma}_{n-1}(z)\eta(z),\;\forall z\in\mathcal{Z}.\end{split}$
Thus, the intractable optimization over the function space $\mathcal{F}_{n-1}$ in Section 4 can be expressed as an optimization over the set of admissible policy profiles $\Pi$ and auxiliary function $\eta(\cdot)$ (see Section C.2 for further details). Note that $\eta(z){}={}\eta(s,\mu,\pi(s,\mu)){}={}\eta(s,\mu)$ for a fixed individual policy $\pi$. This turns $\eta(\cdot)$ into a policy that exerts hallucinated control over the epistemic uncertainty of the confidence set of transitions $\mathcal{F}_{n-1}$ Curi et al. (2020). Furthermore, Equation 8 allows us to optimize over parametrizable functions (e.g., neural networks) $\bm{\pi}$ and $\eta(\cdot)$ using gradient ascent.
We introduce the safety constraint to the objective using the log-barrier method Wright (1992). This restricts the domain on which the objective function is defined only to values that satisfy the constraint Equation 7e, hence, turning Section 4 to an unconstrained optimization problem. Combining these two methods yields the following optimization problem
$\begin{split}\bm{\pi}_{n}^{*}&{}={}\operatorname*{arg\,max}_{\bm{\pi}_{n}\in\Pi}\max_{\eta(\cdot)\in[-1,1]^{p}}\\ \mathbb{E}\Bigg{[}\sum_{t=0}^{T-1}r(\tilde{z}_{n,t})&{}+{}\lambda\log\left(h_{C}(\tilde{\mu}_{n,t{}+{}1})-L_{h}C_{n,t{}+{}1}\right)\Big{\rvert}\tilde{\mu}_{n,0}{}={}\mu_{0}\Bigg{]}\end{split}$ (9a)
subject $\text{subject to}\quad\tilde{a}_{n,t}$ ${}={}\pi_{n,t}(\thinspace_{n,t},\tilde{\mu}_{n,t})$ (9b)
$\tilde{f}_{n-1}(\tilde{z}_{n,t})$ ${}={}\bm{m}_{n-1}(\tilde{z}_{n,t}){}+{}\beta_{n-1}\bm{\Sigma}_{n-1}(\tilde{z}_{n,t})\eta(\tilde{z}_{n,t})$ (9c)
$\thinspace{}_{n,t{}+{}1}$ ${}={}\tilde{f}_{n-1}(\tilde{z}_{n,t}){}+{}\varepsilon_{n,t}$ (9d)
$\tilde{\mu}_{n,t{}+{}1}$ ${}={}U(\tilde{\mu}_{n,t},\pi_{n,t},\tilde{f}_{n-1}),$ (9e)
with $\tilde{z}_{n,t}{}={}(\thinspace_{n,t},\tilde{\mu}_{n,t},\tilde{a}_{n,t})$ and $\lambda>0$ being a tuneable hyperparameter used to balance between the reward and the safety constraint. Provided that the set of safe mean-field distributions (assuming the safe initial distribution $\mu_{0}$ is given) is not empty, $\bm{\pi}_{n}^{*}$ is guaranteed to satisfy the safety constraint during the policy rollout in episode $n$.
Remark 0.
Note that Section 4 can also be used under known transitions by setting $\bm{m}_{n-1}(\cdot)=f(\cdot)$, $\bm{\Sigma}_{n-1}(\cdot){}={}0$ and $L_{h}C_{n,t}=0$, hence, recovering the original constraint $h_{C}(\tilde{\mu}_{n,t}){}\geq{}0$ from Equation 3. In Section 5, we utilize this useful property to construct the upper bound for the reward obtained under unknown transitions.
@@ 4. Safe-M3-UCRL
We summarize the model-based learning protocol used by Safe-M3-UCRL in Algorithm 1. The first step computes constants $C_{n,t}$ (see Appendix A) introduced in Lemma 1. The second step optimizes the objective in Section 4. The third and fourth steps collect trajectories from the representative agent and update the calibrated model. While the learning protocol is model-based, the subroutine in Line 3 can use either model-based or model-free algorithms proposed for the MFC due to our reformulation in Section 4. In Section C.3, we introduce modifications of well-known algorithms for optimizing the mean-field setting.
5. Experiments
In this section, we demonstrate the performance of Safe-M3-UCRL
on the swarm motion benchmark and showcase that it can tackle the real-world large-scale vehicle repositioning problem faced by ride-hailing platforms.
Figure 2. Performance analysis of Safe-M3-UCRL for swarm motion and vehicle repositioning. The policy and statistical model were trained on 10 randomly initialized neural networks for each hyperparameter $p$.
(a) Swarm motion learning curve
(b) Swarm motion mean-field distributions
(c) Swarm motion safety for $p=0.95$
(d) Vehicle repositioning learning curve
(e) Swarm motion policies
(f) Vehicle repositioning safety for $p=0.85$
Performance analysis of the proposed algorithm for swarm motion and vehicle repositioning. The policies and statistical model were trained on 10 randomly initialized neural networks for each hyperparameter $p$ representing the proportion of maximum entropy used as a constraint. The figure displays learning curves, mean-field distributions, safety violation graphs, and policies.
Figure 2. Performance analysis of Safe-M3-UCRL for swarm motion and vehicle repositioning. The policy and statistical model were trained on 10 randomly initialized neural networks for each hyperparameter $p$.
5.1. Swarm Motion
Due to the infancy of Mean-Field RL as a research topic, one of the rare benchmarks used by multiple authors is the swarm motion. Pásztor et al. (2023); Carmona et al. (2019b) view it as Mean-Field RL problem, while Elie et al. (2020) uses it in the context of MFGs. In this setting, an infinite population of agents is moving around toroidal state space with the aim of maximizing a location-dependent reward function while avoiding congested areas Almulla et al. (2017).
Modeling. We model the state space $\mathcal{S}$ as the unit torus on the interval $[0,1]$, and the action space is the interval $\mathcal{A}{}={}[-7,7]$. We approximate the continuous-time swarm motion by partitioning unit time into $T=100$ equal steps of length $\Delta t{}={}1{}/T$. The next state $s_{n,t{}+{}1}{}={}f(z_{n,t}){}+{}\varepsilon_{n,t}$ is induced by the unknown transitions $f(z_{n,t}){}={}s_{n,t}{}+{}a_{n,t}\Delta t$ with $\varepsilon_{n,t}\sim\text{N}(0,\Delta t)$ for all episodes $n$ and steps $t$. The reward function is defined by $r(z_{n,t}){}={}\phi(s_{n,t})-\frac{1}{2}a_{n,t}^{2}-\log(\mu_{n,t})$, where the first term $\phi(s){}={}2\pi^{2}(\sin(2\pi s)-\cos^{2}(2\pi s)){}+{}2\sin(2\pi s)$ determines the positional reward received at the state $s$ (see Section E.3), the second term defines the kinetic energy penalizing large actions, and the last term penalizes overcrowding. Note that the optimal solution for continuous time setting, $\Delta t\rightarrow 0$ can be obtained analytically Almulla et al. (2017) (see Section E.1) and used as a benchmark. Pásztor et al. (2023); Carmona et al. (2019b) show that Mean-Field RL discrete-time, $\Delta t$ 0, methods can learn good approximations of the optimal solution. The disadvantage of these methods is that they can influence the skewness of the mean-field distribution only via overcrowding penalty. Therefore, to control skewness, their only option is to introduce a hyperparameter to the reward that regulates the level of overcrowding penalization. On the other hand, Safe-M3-UCRL controls skewness without trial-and-error reward shaping by imposing the entropic safety constraint $h_{C}(\mu_{n,t}){}={}H(\mu)-C\geq 0$, with $H(\cdot)$ defined in Equation 5, instead of having the overcrowding penalty term $\log(\mu_{n,t})$. Since higher entropy translates into less overcrowding, we can upfront determine and upper-bound the acceptable level of overcrowding by setting a desirable threshold $C$.
@@ 5.1. Swarm Motion
We use a neural network to parametrize the policy profile $\bm{\pi}_{n}=(\pi_{n,0},\ldots,\pi_{n,T-1})$, for every episode $n$, during the optimization of Section 4. The optimization is done by Mean-Field Back-Propagation Through Time (MF-BPTT) (see Section C.3.1). In our experiments, a single neural network shows enough predictive power to represent the whole policy profile, but using $T$ networks, one for each individual policy $\pi_{n,t}$, $t=0,\dots,T-1$, is a natural extension. We use a Probabilistic Neural Network Ensemble Chua et al. (2018); Lakshminarayanan et al. (2017) to represent a statistical model of transitions, which we elaborate in Section C.1. We represent the mean-field distribution by discretizing the state space uniformly and assigning the probability of the representative agent residing within each interval. We set the safety threshold $C$ as a proportion $p\in[0,1]$ of the maximum entropy, i.e., $C{}={}p\max_{\mathcal{P}(\mathcal{S})}H(\mu)$. Note that Safe-M3-UCRL guarantees safe mean-field distributions only if the initial mean-field distributions $\mu_{n,0}$ at time $t=0$ are safe for every episode $n$ for given threshold $C$. A generic way for safe initialization is setting $\mu_{n,0}$ as the maximum entropy distribution among all safe distributions $\zeta$
$\mu_{n,0}=\operatorname*{arg\,max}_{\mu\in\zeta}H(\mu).$ (10)
Note that, in general, the safe initial distribution might not exist. We provide the implementation details in Section D.2.
Results. In Figure 1(a), we observe the learning curve of Safe-M3-UCRL for $p=0.95$ for 10 randomly initialized runs. The learning process is volatile in the initial phase due to the high epistemic uncertainty, but after 50 episodes, all policies converge toward the solution as if the transitions were known. In Figure 1(b), we use various thresholds $C$ to show that the entropic constraint effectively influences the degree of agents’ greediness to collect the highest positional reward. By increasing $p$ towards 1, we force agents to put increasingly more emphasis on global welfare rather than on individual rewards. We see that for $p=0.5$ we obtain a distribution that matches the distribution obtained by the unconstrained M3-UCRL  Pásztor et al. (2023) that relies on the overcrowding penalty, while for $p=0.95$ we significantly surpass the effect that the penalty has on the distribution’s skewness. We also show that the discrete-time solutions with low $p$ serve as a good approximation of the continuous-time optimal distribution $\mu^{*}$. Furthermore, we observe that the policies under unknown transitions overlap with the solutions learned under known transitions. In Figure 1(e), we observe that unconstrained policies and policies with low $p$ push agents towards high individual rewards. Note that for states close to 1, the algorithms learn it requires less kinetic energy to push agents over the border due to the toroidal shape of the state space. For high $p$’s learned policies push agents always in the same direction to maintain uniformity of the system. Importantly, Figure 1(c) shows that Safe-M3-UCRL for $p=0.95$ keeps the mean-field distribution safe throughout the entire execution, unlike the solutions that rely on the overcrowding penalty term. For further details, see Section E.3.
Figure 3. Safe-M3-UCRL guided vehicle distribution in Shenzhen in the evening peak hours.
Observed demand $\rho_{0}$ used as a target distribution
Unconstrained M3-UCRL under known transitions
Safe-M3-UCRL under known transitions for $p=0.85$
Safe-M3-UCRL under unknown transitions for $p=0.85$
The proposed algorithm guided vehicle distribution in Shenzhen in the evening peak hours. The image shows heatmaps illustrating the spatial distribution of vehicles.
Figure 3. Safe-M3-UCRL guided vehicle distribution in Shenzhen in the evening peak hours.
5.2. Vehicle Repositioning Problem
Since ride-hailing services, such as Uber, Lyft, and Bolt, gained popularity and market share, vehicle repositioning has been a long-standing challenge for these platforms, i.e., moving idle vehicles to areas with high-demand potential. A similar challenge is present in bike-sharing services accessible in many cities and, as of more recently, dockless electric scooter-sharing services such as Bird and Lime. In the competitive environment, the operator significantly increases profit by successfully repositioning idle vehicles to high-demand areas. Nevertheless, there might exist regulations imposed by the countries or cities that enforce service providers to either guarantee fair service accessibility or restrict the number of vehicles in districts with high traffic density. Such restrictions prevent operators from greedily maximizing the profit and can be encapsulated by some dispersion metric such as entropy. Solving this problem helps prevent prolonged vehicle cruising and extensive passenger waiting times in the demand hotspots, increasing the service provider’s efficiency and reducing its carbon footprint. Existing approaches to vehicle repositioning range from static optimization over a queuing network Zhang and Pavone (2016); Braverman et al. (2019), model predictive control Iglesias et al. (2018), to RL Wen et al. (2017); Lin et al. (2018); Mao et al. (2020). The main advantage of Safe-M3-UCRL is the capability of controlling a large fleet of homogeneous vehicles and enforcing efficient coordination to match the spatiotemporal demand distribution. Additionally, the safety constraint introduced into the model guarantees service accessibility by ensuring idle vehicles are spreading over the study region. Although accessibility has not been widely discussed in the literature on vehicle repositioning, it is expected to be an important fairness constraint when shared mobility services become a prevailing travel mode Shaheen et al. (2017).
Modeling. Ride-hailing operations can be modeled as sequential decision-making, which consists of passenger trips followed by repositioning trips operated by a central controller as illustrated in Figure 1. We assume that the controller has access to the locations of vehicles in its fleet and communicates the real-time repositioning actions to the drivers via electronic devices. Nevertheless, since the fleet is operating in a noisy traffic environment, repositioning usually cannot be executed perfectly. We assume that vehicles can move freely within the area of our interest, which is represented by a two-dimensional unit square, i.e., the state-space $\mathcal{S}{}={}[0,1]^{2}$, and repositioning actions are taken from $\mathcal{A}{}={}[-1,1]^{2}$. The objective of our model is to satisfy the demand in the central district of Shenzhen while providing service accessibility in the wider city center. We restrict our modeling horizon to three evening peak hours, which are discretized in fifteen-minute operational intervals, i.e., $T=12$, and each episode $n$ represents one day (for more details, see Appendix D). We model service providers goal of maximizing the coverage of the demand by the negative of the Kullback-Leibler divergence between the vehicles’ distribution $\mu_{n,t}$ and demand for service denoted as $\rho_{0}$, i.e., $r(z_{n,t}){}={}-D_{KL}(\rho_{0}||\mu_{n,t})$. In particular, the demand distribution $\rho_{0}\in\mathcal{P}(\mathcal{S})$ represents a probability of a trip originating in the infinitesimal neighborhood of state $s\in\mathcal{S}$ during peak hours (see Figure 2(a)). We estimate a stationary demand distribution $\rho_{0}$ from the vehicle trajectories collected in Shenzhen, China, in 2016. If the passenger’s trip originates at $s\in\mathcal{S}$ the likelihood of its destinations is defined by the mapping $\Phi:\mathcal{S}\rightarrow\mathcal{P}(\mathcal{S})$, which we fit from the trip trajectories (see Section D.1). We use $\Phi(\cdot)$ to define sequential transitions by first executing passenger trips followed by vehicle repositioning. Formally, the next state $s_{n,t{}+{}1}{}={}f(z_{n,t}){}+{}\varepsilon_{n,t}$ is induced by the unknown transitions $f(z_{n,t}){}={}\text{clip}(s^{\Phi}_{n,t}{}+{}a_{n,t},0,1)$, where $s^{\Phi}_{n,t}\sim\Phi(s_{n,t})$ and $\varepsilon_{n,t}\sim\text{TN}(0,\sigma^{2}I_{2})$ is a Gaussian with a known variance $\sigma^{2}$ truncated at the borders of $\mathcal{S}$ and $I_{2}$ is the $2\times 2$ unit matrix. Notice that the controller determines repositioning actions given intermediate states $s^{\Phi}_{n,t}$ obtained after executing passenger trips. We use entropic safety constraint $h_{C}(\mu_{n,t}){}={}H(\mu)-C\geq 0$ to enforce the service accessibility across all residential areas (see Figures 2(c) to 2(d)). Therefore, the optimization objective in Section 4 trades off between greedily satisfying the demand $\rho_{0}$ and adhering to accessibility constraint imposed by $h_{C}(\cdot)$. Identically to the swarm motion experiment, we use a neural network to parametrize the policy profile $\bm{\pi}_{n}$, which we optimize by MF-BPTT. A statistical model of the transitions is represented by a Probabilistic Neural Network Ensemble, while $\mu_{n,0}$ is initialized using Equation 10. We represent the mean-field distribution by discretizing the state space into the uniform grid as elaborated in Section D.1.